400 likes | 716 Views
Tracker Summary. Makoto Yoshida Osaka Univ. MICE CM @ Frascati June 28 th , 2005. Talks in tracker session. Emittance measurement simulation update (C. Rogers) Tracker software status (M. Ellis / A. Bross) Tracker Status (G. Barber / K. Long) Tracker electronics update (A. Bross)
E N D
Tracker Summary Makoto Yoshida Osaka Univ. MICE CM @ Frascati June 28th, 2005
Talks in tracker session • Emittance measurement simulation update (C. Rogers) • Tracker software status (M. Ellis / A. Bross) • Tracker Status (G. Barber / K. Long) • Tracker electronics update (A. Bross) • Tracker DAQ (M.Yoshida) • KEK beam test strategy (K. Yoshimura) • KEK beam test analysis (M. Yoshida)
Tracker Software Status M. Ellis MICE Collaboration Meeting 27th June 2005
Progress so far • G4MICE has been extended to handle calibration, alignment and decoding information. Further work is needed before it will be ready to be committed. • Visualisation framework has been developed (based on the libsx wrapper of X) with a simple test program that can be extended to handle both test-beam and simulation needs. • Pattern recognition and track fitting code modified to work with no magnetic field.
Plans • Due to lack of effort available, G4MICE activities will take a second priority to test-beam readout until it is finished. • At that stage, work described so far will be completed, hopefully in time to be used for monitoring and analysis of test-beam data. • If time is available, still aim to study optimum station spacing before next collaboration meeting.
Tracker Status • Contents:- • Station Fabrication • Modifications to the Existing Stations • Mechanical Support • Station Connectors • Complete Prototype • Changes for Final Tracker
Complete Prototype(Prior to outer covers being fitted) Status – The completed prototype was shipped to FermiLab 20 June for cosmic ray tests and will then be forwarded to KEK for the beam tests
Stephanie Wang, Wing Lau and Geoff met to discuss space at the end of the tracker Issues were: Support of diffuser mechanism Avoid compromising integrity of vacuum-vessel seal Support of iron sheild Avoid possibility that support for shield will interfere with patch panel Patch panel, diffuser support and iron shield
The stations will have a thickened pad around the 6 fixation points. This will allow for machining to ensure a uniform thickness. A quality control regimen needs to be in place for each stage of construction/assembly Waveguide patch panel connectors will need to be ‘turned around’ The location/relocation mechanism in the bore of the cryostat will need to be designed and tested The Patch panel will need the final design and pressure calculations completed Vacuum test of station assembly to be performed over the summer Plan to construct a model of complete device to ensure no space conflicts during manufacture or installation Changes for Final Tracker
MICE Tracker Electronics Update Progress towards a tracker test at KEK
Status: VLPC System • The VLPC system is complete and operating to spec • Can easily control to the 9K spec. • With heaters off T=5.3K • Uses a Sumitomo SRDK-415D Cryo-cooler. • This is the unit we plan to use throughout MICE
Status: Cryo System • Full Commissioning done – Performance To Spec!!!
Status:Electronics • The system will use 4 AFE II boards from D0. These are now working 512 ch ADC Discrim. TDC in next version
Status Electronics • 6 AFE II boards have now been fully tested and are available to MICE • All Channels tested • Temperature and bias control operation • LVDS readout functional • Still need to run the boards through the temperature and bias control calibration test stands
Timescale • Next two weeks • Get 4 board readout working • Get sequencer controller replacement working • July 18-22 • Calibrate VLPC Cassettes • July 25-29 • Cosmic Ray Test • August 1 • Pack and Ship to KEK • August 8-12 • Preliminary setup and shakeout at KEK • At least operate VLPC cryo 大丈夫
Tracker DAQ Makoto Yoshida (Osaka Univ.) MICE meeting @ Frascati 2005/6/27
Status • Setup LinuxPC with Bit3 VME interface at Fermilab in Dec. 2005 • Start porting the Windows VB routines to C codes on Linux PC • Successful readout pedestal data on a AFEII board in the end of May. • Initialization AFE boards via 1553 • Control SASeq • Readout buffer module, VLSB • GUI is prepared for the test • VME ADC and TDC modules have been tested in May KEK beam test to readout TOF, ACC and other beam counters in UniDAQ framework
Tracker DAQ at Fermilab Backplane AFEII Cryocooler Bit3 SASeq 1553 VLSB
To do for KEK beam test • Extend slow-controller to handle 4 boards of AFEII • Implement AFEII readout routines on UniDAQ framework • Merge AFEII data with the other ADC/TDC data • Integration test at Fermilab • Time stamp of integration window on Trip chip • Tracker prototype will be shipped from UK in July • LED and Cosmic-ray data taking • Setup the system in KEK in Aug.
To do for MICE • Event-buffering in VLSB • Need to trigger 1MHz (1k muons in 1msec from ISIS) • Implement zero suppression scheme • Bench mark test of readout speed • Establish event numbering scheme / synchronization to the other counters
KEK beam test • Check the performance of prototype tracker at KEK-PS p2 beam line • Step 0 • Feb. 2005 • Step 1 • May 2005 • Step 2 • Sep. 2005 (not yet approved)
Q3 Q4 D4 Setup T1 2cm-thick scintillator Aerogel Diffuser Defining Counters TOF hodoscope
Timing resolutionw/ and w/o magnetic field 3GeV/c p- Solenoid OFF Solenoid ON • = 77ps • TOFcounter=65ps • = 80ps • TOFcounter=68ps After time walk correction The system can measure time of flight with 80ps resolution even in the fringe magnetic field need more offline work to obtain better resolution than 70ps
ACC performance 0.425GeV/c Negative e- m- p-
Diffuser study -0.3GeV/c Beam divergence w/o solenoid increase to ~20cm/200cm=100mrad @15mmPb
Muon yield in low momentum beam TOF distributions -0.3GeV/c 5475 m / 16min =6Hz e- p- m- -0.4GeV/c 9000 m / 10min =15Hz
Summary of Step 1 (May2005) • Beam test with SC solenoid was performed at KEK-PS p2 beam line during May 26th– June 2nd • TOF performance was checked in magnetic field • Intrinsic resolution of the counter better then 70ps • Resolution of the TOF system ~ 80ps • Need more careful offline work • New ACC (refractive index = 1.05) was tested • good performance to discriminate p/m/e at 0.425GeV/c • Beam parameter was surveyed to be optimized for Autumn beam test • The effect of lead diffuser was measured, will be compared with simulations. • VME ADC/TDC modules were tested to take beam data • Ready to test tracker prototype in autumn Step 2